OpenStudy (anonymous):

solve for x: 6.05 = 8+x-3x^3

OpenStudy (muzzack):

1. Swap Both Sides of Equation 8+x−3*x^3=6.05 2. Negate 8+x−3*x^3=6.05 3. Multiply Terms 8+x−3x^3=6.05 4. Order Terms -3x^3+x+8=6.05 5. Convert Decimals to Fractions -3x^3+x+8=605/100 6. Eliminate Greatest Common Denominator -3x^3+x+8=121/20 7. Move Constants to Right -3x^3+x=121/20−8 8. Add Fractions -3x^3+x=-8*20+121/20 9. Multiply Terms -3x^3+x=-160+121/20 10. Add Constants -3x^3+x=-39/20 11. Convert Improper Fractions to Proper Fractions -3x^3+x=-119/20

OpenStudy (anonymous):

is there more?